# Vulnerabilities of smart contracts and mitigation schemes: A Comprehensive Survey

### v1

@@ IV-B Dynamic Analysis
The term ”dynamic analysis” refers to the observation of code while it is being executed in its original context. It is also known as black-box testing because it is performed without access to the source code. It works in the same way as an attacker who feeds malicious code or unpredictable input to the appropriate functions of a program to look for vulnerabilities. The most common technique is Fuzzing.
IV-B1 Fuzzing
@@
Fuzzy testing, often known as fuzzing, is a kind of automated software testing that involves injecting incorrect, malformed, or unpredictable inputs into a system; the objective is to uncover software flaws and vulnerabilities. A fuzzing tool injects these inputs into the program and then watches for problems such as crashes or data leaks; it also can perform static analysis on execution traces. In the case of smart contracts, Wang et al. [53] deploy the smart contract on a test network; then, they monitor the balance of the smart contract to identify basic misappropriation. This technique removes the necessity for particular patterns to identify a vulnerability.
IV-C Formal Verification
@@
Formal verification is a technique that automates bug detection of a hardware or software system by comparing a formal model of the system to formal requirements or behaviour specifications. Through mathematical analysis, formal verification can provide a high level of confidence.
@@
In order to conduct a formal verification, first we need to provide a specification. A program specification is an unambiguous definition the purpose of the program as well as the scenarios that are allowed or not to be executed. In order to verify a property, tools usually build deduction trees where the root presents a Hoare triplet [54]. A Hoare triplet consists of: (a) Pre condition: It is the initial state; (b) Instructions: A series of transactions; and (c) Post condition: It is the property to verify. In order to build deduction trees theorem proving algorithms are usually used such as Coq [55] and Isabelle/HOL [54].
@@
There are a few contributions that propose languages to write formal specifications. Permenev et al. [56] proposed a specification language, called VerX, where the syntax is similar to Solidity; it supports temporal logic. The Ethereum community also proposed a language named Act [57] which is a specification for Ethereum Virtual Machine (EVM) programs.
IV-D Comparison
@@
The goal of using detection methods such as static, dynamic, and formal verification is to ensure that smart contracts are correctly executed in all states; this is to ensure that no vulnerabilities are produced to ensure that specifications are respected. Table V shows the papers that discuss each methodology.
@@
Static code analysis tools can potentially produce false negatives results; vulnerabilities occur but are not reported. This could happen because the analysis tool used doesn’t allow for a large number of test scenarios; it doesn’t go into detail when looking at different states because it usually gives you a timeout. Various vulnerabilities may be false negatives in static analysis; however, they can be detected correctly utilizing dynamic analysis tools. Thus, it is strongly recommended to combine both types of analysis.
@@
Dynamic analysis generates transactions with random inputs. This means that it will observe random states depending on inputs. The challenge for these techniques is to generate enough transactions and to carefully select inputs to achieve maximum coverage.
@@
Formal verification can provide full coverage for the specification, under consideration, because it employs mathematical analysis. This guarantees the satisfaction of a property if verified; however, in general, we cannot cover everything because the verification can become easily intractable. Thus, false negatives can still be present. In addition, formal verification requires a skilled programmer who can express a smart contract as a mathematical high-level specification while accounting for a specific low-level virtual machine. This operation takes a long time and requires a lot of resources. Audit firms usually make use of formal verification [58].
V Vulnerability Detection Tools
@@
In this section, we will cover the most popular tools for detecting smart contract vulnerabilities. To investigate available tools, we first looked through academic literature and review articles (see Table VI). Then we covered five of the most commonly used tools as indicated by developer forums as well as the number of forks and update frequency on GitHub.
# | Tool | Detection Technique | Last update |
1 | Slither [6] [67] | Static Analysis | May 2023 |
2 | MythX [68] | static, dynamic & Symbolic Execution | Not Public |
3 | Mythril [7] [69] | Symbolic Execution | May 2023 |
4 | Echidna [9] [70] | Fuzzer | Apr 2023 |
5 | Manticore [8] [71] | Symbolic Execution | June 2022 |
6 | Securify [39] [72] | Static Analysis | Sep 2021 |
7 | KEVM [60] [73] | Static Analysis | Apr 2022 |
8 | Smartcheck [18] [74] | Static Analysis | Dec 2019 deprecated |
9 | MadMax [42] [75] | static analysis | Jun 2021 |
10 | Vertigo [76] [77] | Mutation Testing | Feb 2021 |
11 | EtherSolve [38] [78] | CFG Extraction | Nov 2021 |
12 | Octopus [79] | Symbolic Execution | Nov 2020 |
13 | Oyente [5] [80] | Symbolic Execution | Nov 2020 deprecated |
14 | ERC20 Verifier [81] | Verify ERC20 Compatibility | Nov 2019 |
15 | Solgraph [82] | CFG Extraction | Jan 2019 |
16 | Osiris [41] [83] | Symbolic Execution | Sep 2018 |
V-A Oyente
V-A1 Description
@@
Luu et al. propose Oyente [5] the first symbolic execution tool for Ethereum smart contracts. Oyente was the basis of several tools developed later [62] [41]. It has four components: (a) CFGBuilder: It constructs CFG of the smart contract; (b) Explorer: It takes as input the Ethereum state. It has a loop that runs a state then executes an instruction on the output of that state. The loop continues until no state remains or a timeout is reached. The output is a symbolic trace. Explorer determines the infeasible trace by querying the Z3 SMT solver [84]; (c) CoreAnalysis: It targets predefined vulnerabilities by looking for patterns on the symbolic trace; (d) Validator: It queries Z3 solver with traces flagged as vulnerable to reduce false positives cases.
@@
Oyente detects seven types of vulnerabilities, namely Re-entrancy, Integer overflow/underflow, Transaction order dependence, Timestamp dependence, Callstack Depth, EVM Code Coverage, and Parity Multisig bugs.
V-A2 Execution Example
@@
The primary aim is to assess Oyente’s effectiveness in detecting vulnerabilities that could also impact contracts handling fractionalized NFTs. The results, depicted in Figure 17, highlight an Integer Overflow vulnerability, along with the execution trace where this issue is detected.
V-A3 Pros and Cons
Oyente’s deployment via a Docker image significantly simplifies its setup process, making it accessible for users with varying levels of expertise in blockchain technology. This ease of use facilitates quick integration into development workflows, allowing for the immediate analysis of Ethereum smart contracts. Despite its user-friendly setup, Oyente’s compatibility is somewhat limited; it supports Ethereum compiler versions only up to 0.4.17. This restriction may hinder its applicability to smart contracts compiled with newer versions of the Solidity compiler, potentially limiting its utility in analyzing the latest smart contract developments.
The output provided by Oyente is both clear and concise, enabling developers and auditors to easily interpret vulnerability reports. Moreover, Oyente has played a pivotal role in the evolution of smart contract analysis tools. It has laid the groundwork for subsequent innovations, including Maian [86] and Osiris [41]. These tools have built upon Oyente’s foundational techniques, such as the construction of Control Flow Graphs (CFGs), to offer enhanced accuracy in detecting vulnerabilities and orphan paths within smart contracts.
V-B Slither
V-B1 Description
@@
Slither is proficient in uncovering 70 bug varieties, which encompass self-destructive smart contracts, code injections via delegatecall, frozen ether, and Reentrancy vulnerabilities [87].
V-B2 Execution Example
We analyzed the FractionalNFTMarket smart contract in Listing 18 using Slither to identify potential security vulnerabilities. The results of the analysis are presented in Figure 19. Slither has successfully detected a reentrancy vulnerability within the sellShares function of the contract, which could allow an attacker to exploit the contract by recursively calling the function to drain its funds. In addition to detecting this critical vulnerability, Slither offered recommendations for code improvements and provided a comprehensive set of information, including a human-readable summary, an inheritance graph, and the Control Flow Graph (CFG) of each function.
Slither also provides a list of useful information including a human readable summary of the scanned contract, inheritance graph, and CFG of each function; the complete list can be found in the github repository [67].
V-B3 Pros and Cons
@@
One limitation of Slither is its tendency to report false positives, which can complicate the auditing process by requiring additional validation to go through the findings. For a more comprehensive analysis, Crytic [89], a premium service, can be utilized. Crytic extends Slither’s capabilities by identifying an additional 50 types of faults that Slither might not catch. Furthermore, Crytic offers integration with GitHub, allowing for automated testing on pull requests, thereby enhancing the development workflow and ensuring continuous contract integrity.
V-C Mythril
V-C1 Description
@@
Slithe Durieux et al. [7] developed a tool known as Mythril that is capable of identifying potential weaknesses in smart contracts by utilizing symbolic execution. Mythril, a Python-based command-line tool, allows for interactive inspection of smart contracts. It produces a Control Flow Graph (CFG) and performs symbolic execution of EVM bytecode to restrict the search area. The tool can generate concrete values to exploit any detected vulnerabilities. To evaluate the feasibility of different paths, Mythril makes use of the Z3 SMT solver [84].
@@
Mythril has the ability to discover fourteen distinct types of vulnerabilities including Delegate Call To Untrusted Contract, Dependence on Predictable Variables, Deprecated Opcodes, Ether Thief, Exceptions, External Calls, Integer Over/Underflow, Multiple Sends, Self-Destruction, State Change External Calls, Unchecked Return Values, User Supplied Assertions, and Arbitrary Storage Write and Arbitrary Jump [90].
V-C2 Execution Example
@@
The FractionalNFTMarket.sol smart contract, presented in Listing 18, has been analyzed using Mythril to uncover potential security issues. The resulting output is depicted in Figure 20. Mythril has identified a reentrancy vulnerability within the contract’s ‘sellShares‘ function. The output provides detailed information, including the state of the contract when the vulnerability can be triggered and the specific transaction sequence that leads to the issue. Unlike Slither, Mythril offers not only the identification of vulnerabilities but also the concrete inputs that cause them, enhancing the developer’s understanding of the contract’s weaknesses. Nevertheless, Mythril’s output does not extend to include recommendations for remediation of the detected issues.
V-C3 Pros and Cons
@@
Mythril stands out for its accessibility, offering straightforward installation options via Docker or Python package manager, which streamlines the setup process. It allows for granular analysis, with the flexibility to test individual functions through custom Python scripts. This targeted testing can be particularly advantageous for developers seeking to debug specific aspects of their smart contract code.
@@
However, Mythril’s thorough approach to smart contract analysis can be computationally intensive, often resulting in longer analysis times compared to tools like Slither. The resource-heavy nature of Mythril’s symbolic execution process means that it may not be the most efficient choice for rapid iteration during development. For those seeking a more advanced feature set and cloud-based capabilities, Mythril’s commercial counterpart, MythX [68], offers enhanced analysis power and can be seamlessly integrated into continuous integration/continuous deployment (CI/CD) pipelines [91].
V-D Manticore
V-D1 Description
@@
Mossberg et al. [8] implemented a tool named Manticore to detect vulnerabilities in smart contracts. The primary aim of Manticore is to track inputs that kill a program, record instruction-level implementation, and provide Python API access to its analysis engine. To locate distinct computation paths in EVM bytecode, Manticore uses symbolic execution. It discovers inputs that will stimulate these computation paths with the aid of the SMT solver Z3 [84]. It keeps track of the execution traces for each execution. Manticore converts Solidity code to bytecode for evaluation; then, it examines the traces for vulnerabilities, such as reentrancy and reachable self destruct operations, reporting them in the context of the source code. This tool is developed by TrailOfBits [92].
@@
Mythril detects ten types of vulnerabilities namely, delegatecall, overflow, reentrancy, Use of potentially unsafe/manipulable instructions, Reachable external, Reachable selfdestruct instructions, Uninitialized memory usage, Uninitialized storage usage, Enable INVALID instruction detection, and Unused internal transaction return values [93].
V-D2 Execution Example
@@
For our examination of Manticore, we used the TokenSaleChallenge.sol [85] same one used to test Oyente. This contract represents a simplified model for trading tokens, which can be seen as stand-ins for fractional tokens of an NFT.
@@
The tool is extremely slow as it takes more than one hour to perform the test; it takes less than a minute for the other tools using the same smart contract and machine. The output of the analysis is a folder containing a summary report as well as traces and analysis for each test case. Manticore generated 51 test-cases, but it failed to detect the smart contact’s reentrancy vulnerability, as shown in Figure 21.
V-D3 Pros and Cons
@@
Manticore offers a diverse set of command-line tools along with scriptable Python APIs, which is suitable to a variety of use cases and allow for in-depth analysis procedures. Manticore’s cover a broad spectrum of known vulnerabilities, making it a valuable asset for developers seeking to secure their smart contracts.
@@
However, the detailed nature of Manticore’s analysis comes with a trade-off in terms of performance. The tool can take significantly longer to complete its analysis compared to other tools, sometimes leading to timeouts. This is often due to its comprehensive symbolic execution approach, which, while powerful, is also resource-intensive. Additionally, Manticore is known to require a considerable amount of memory, which could be a limiting factor for users with constrained hardware resources.
V-E Echidna
V-E1 Description

### v2

@@ 4.2 Dynamic Analysis
”dynamic analysis” refers to observing code while executed in its original context. It is also known as black-box testing because it is performed without access to the source code. It works the same way as an attacker who feeds malicious code or unpredictable input to the appropriate functions of a program to look for vulnerabilities. The most common technique is Fuzzing.
4.2.1 Fuzzing
@@
Fuzzy testing, often known as fuzzing, is a kind of automated software testing that involves injecting incorrect, malformed, or unpredictable inputs into a system; the objective is to uncover software flaws and vulnerabilities. A fuzzing tool injects these inputs into the program and then watches for problems such as crashes or data leaks; it also can perform static analysis on execution traces. In the case of smart contracts, Wang et al. [53] deploy the smart contract on a test network; then, they monitor the balance of the smart contract to identify basic misappropriation. This technique removes the necessity for particular patterns to determine a vulnerability.
4.3 Formal Verification
@@
Formal verification is a technique that automates bug detection of a hardware or software system by comparing a formal system model to formal requirements or behavior specifications. Through mathematical analysis, formal verification can provide a high level of confidence.
@@
To conduct a formal verification, we first need to provide specifications. A program specification is an unambiguous definition of the program’s purpose and the scenarios that are allowed or not to be executed. Tools usually build deduction trees to verify a property where the root presents a Hoare triplet [54]. A Hoare triplet consists of: (a) Pre condition: It is the initial state; (b) Instructions: A series of transactions; and (c) Post condition: It is the property to verify. To build deduction trees, theorem-proving algorithms are usually used, such as Coq [55] and Isabelle/HOL [54].
@@
There are a few contributions that propose languages to write formal specifications. Permenev et al. [56] proposed a specification language called VerX, where the syntax is similar to Solidity; it supports temporal logic. The Ethereum community also proposed a language named Act [57], which is a specification for Ethereum Virtual Machine (EVM) programs.
4.4 Comparison
@@
The goal of using detection methods such as static, dynamic, and formal verification is to ensure that smart contracts are correctly executed in all states; this is to ensure that no vulnerabilities are produced and that specifications are respected. Table V shows the papers that discuss each methodology.
@@
Static code analysis tools can potentially produce false negative results; vulnerabilities occur but are not reported. This could happen because the analysis tool doesn’t allow for many test scenarios; it doesn’t go into detail when looking at different states because it usually gives you a timeout. Various vulnerabilities may be false negatives in static analysis; however, they can be detected correctly utilizing dynamic analysis tools. Thus, it is strongly recommended that both types of analysis be combined.
@@
Dynamic analysis generates transactions with random inputs, which means that it observes random states depending on the inputs. The challenge for these techniques is to generate enough transactions and carefully select inputs to achieve maximum coverage.
@@
Formal verification can provide full coverage for the specification under consideration because it employs mathematical analysis. This guarantees the satisfaction of a property if verified; however, in general, we cannot cover everything because the verification can become quickly intractable. Thus, false negatives can still be present. In addition, formal verification requires a skilled programmer who can express a smart contract as a mathematical high-level specification while accounting for a specific low-level virtual machine. This operation takes a long time and requires a lot of resources. Audit firms usually make use of formal verification [58].
5 Vulnerability Detection Tools
@@
This section will cover the most popular tools for detecting smart contract vulnerabilities. To investigate available tools, we first looked through academic literature and review articles (see Table VI). Then, we covered five of the most commonly used tools, as indicated by developer forums, and the number of forks and update frequency on GitHub.
Tool |
Detection Technique |
Last update |
1 |
Static Analysis |
May 2023 |
2 |
MythX [68] |
static, dynamic & Symbolic Execution |
Not Public |
3 |
Symbolic Execution |
May 2023 |
4 |
Fuzzer |
Apr 2023 |
5 |
Symbolic Execution |
June 2022 |
6 |
Static Analysis |
Sep 2021 |
7 |
Static Analysis |
Apr 2022 |
8 |
Static Analysis |
Dec 2019 deprecated |
9 |
static analysis |
Jun 2021 |
10 |
Mutation Testing |
Feb 2021 |
11 |
CFG Extraction |
Nov 2021 |
12 |
Octopus [79] |
Symbolic Execution |
Nov 2020 |
13 |
Symbolic Execution |
Nov 2020 deprecated |
14 |
ERC20 Verifier [81] |
Verify ERC20 Compatibility |
Nov 2019 |
15 |
Solgraph [82] |
CFG Extraction |
Jan 2019 |
16 |
Symbolic Execution |
Sep 2018 |
5.1 Oyente
5.1.1 Description
@@
Luu et al. propose Oyente [5], the first symbolic execution tool for Ethereum smart contracts. Oyente was the basis of several tools developed later [62] [41]. It has four components: (a) CFGBuilder: It constructs CFG of the smart contract; (b) Explorer: It takes as input the Ethereum state. It has a loop that runs a state and then executes an instruction on the output of that state. The loop continues until no state remains or a timeout is reached. The output is a symbolic trace. Explorer determines the infeasible trace by querying the Z3 SMT solver [84]; (c) CoreAnalysis: It targets predefined vulnerabilities by looking for patterns on the symbolic trace; (d) Validator: It queries Z3 solver with traces flagged as vulnerable to reduce false positives cases.
@@
Oyente detects seven types of vulnerabilities: Re-entrancy, Integer overflow/underflow, Transaction order dependence, Timestamp dependence, Callstack Depth, EVM Code Coverage, and Parity Multisig bugs.
5.1.2 Execution Example
@@
The primary aim is to assess Oyente’s effectiveness in detecting vulnerabilities that could impact contracts handling fractionalized NFTs. The results, depicted in Figure 17, highlight an Integer Overflow vulnerability and the execution trace where this issue is detected.
5.1.3 Pros and Cons
Oyente’s deployment via a Docker image significantly simplifies its setup process, making it accessible for users with varying expertise in blockchain technology. This ease of use facilitates quick integration into development workflows, allowing for the immediate analysis of Ethereum smart contracts. Despite its user-friendly setup, Oyente’s compatibility is somewhat limited; it supports Ethereum compiler versions only up to 0.4.17. This restriction may hinder its applicability to smart contracts compiled with newer versions of the Solidity compiler, potentially limiting its utility in analyzing the latest smart contract developments.
The output provided by Oyente is clear and concise, enabling developers and auditors to interpret vulnerability reports easily. Moreover, Oyente has played a pivotal role in evolving smart contract analysis tools. It has laid the groundwork for subsequent innovations, including Maian [86] and Osiris [41]. These tools have built upon Oyente’s foundational techniques, such as constructing Control Flow Graphs (CFGs), to offer enhanced accuracy in detecting vulnerabilities and orphan paths within smart contracts.
5.2 Slither
5.2.1 Description
@@
Slither is proficient in uncovering 70 bug varieties, encompassing self-destructive smart contracts, code injection via delegatecall, frozen ether, and Reentrancy vulnerabilities [87].
5.2.2 Execution Example
Using Slither to identify potential security vulnerabilities, we analyzed the FractionalNFTMarket smart contract in Listing 18. The analysis results are presented in Figure 19. Slither has successfully detected a reentrancy vulnerability within the sellShares function, which could allow an attacker to exploit the contract by recursively calling the function to drain its funds. In addition to detecting this critical vulnerability, Slither offered recommendations for code improvements and provided a comprehensive set of information, including a human-readable summary, an inheritance graph, and the Control Flow Graph (CFG) of each function.
Slither also provides a list of helpful information, including a human-readable summary of the scanned contract, inheritance graph, and CFG of each function; the complete list can be found in the GitHub repository [67].
5.2.3 Pros and Cons
@@
One limitation of Slither is its tendency to report false positives, which can complicate the auditing process by requiring additional validation to review the findings. Crytic [89], a premium service, can be utilized for a more comprehensive analysis. Crytic extends Slither’s capabilities by identifying an additional 50 types of faults that Slither might miss. Furthermore, Crytic offers integration with GitHub, allowing for automated testing on pull requests, thereby enhancing the development workflow and ensuring continuous contract integrity.
5.3 Mythril
5.3.1 Description
@@
Slithe Durieux et al. [7] developed a tool known as Mythril to identify potential weaknesses in smart contracts by utilizing symbolic execution. Mythril, a Python-based command-line tool, allows for interactive inspection of smart contracts. It produces a Control Flow Graph (CFG) and performs symbolic execution of EVM bytecode to restrict the search area. The tool can generate concrete values to exploit any detected vulnerabilities. To evaluate the feasibility of different paths, Mythril uses the Z3 SMT solver [84].
@@
Mythril can discover fourteen distinct types of vulnerabilities, including Delegate Call To Untrusted Contracts, Dependence on Predictable Variables, Deprecated Opcodes, Ether Thief, Exceptions, External Calls, Integer Over/Underflow, Multiple Sends, Self-Destruction, State Change External Calls, Unchecked Return Values, User Supplied Assertions, and Arbitrary Storage Write and Arbitrary Jump [90].
5.3.2 Execution Example
@@
The FractionalNFTMarket.sol smart contract, presented in Listing 18, has been analyzed using Mythril to uncover potential security issues. The resulting output is depicted in Figure 20. Mythril has identified a reentrancy vulnerability within the contract’s ‘sellShares‘ function. The output provides detailed information, including the state of the contract, when the vulnerability can be triggered, and the specific transaction sequence that leads to the issue. Unlike Slither, Mythril offers the identification of vulnerabilities and the concrete inputs that cause them, enhancing the developer’s understanding of the contract’s weaknesses. Nevertheless, Mythril’s output does not extend to include recommendations for remediation of the detected issues.
5.3.3 Pros and Cons
@@
Mythril stands out for its accessibility, offering straightforward installation options via Docker or Python package manager, streamlining the setup process. It allows for granular analysis, with the flexibility to test individual functions through custom Python scripts. This targeted testing can be particularly advantageous for developers seeking to debug specific aspects of their smart contract code.
@@
However, Mythril’s thorough approach to smart contract analysis can be computationally intensive, often resulting in longer analysis times than tools like Slither. The resource-heavy nature of Mythril’s symbolic execution process may not be the most efficient choice for rapid iteration during development. For those seeking a more advanced feature set and cloud-based capabilities, Mythril’s commercial counterpart, MythX [68], offers enhanced analysis power and can be seamlessly integrated into continuous integration/continuous deployment (CI/CD) pipelines [91].
5.4 Manticore
5.4.1 Description
@@
Mossberg et al. [8] implemented a tool named Manticore to detect vulnerabilities in smart contracts. The primary aim of Manticore is to track inputs that kill a program, record instruction-level implementation, and provide Python API access to its analysis engine. Manticore uses symbolic execution to locate distinct computation paths in EVM bytecode. It discovers inputs that will stimulate these computation paths with the aid of the SMT solver Z3 [84]. It keeps track of the execution traces for each execution. Manticore converts Solidity code to bytecode for evaluation; then, it examines the traces for vulnerabilities, such as reentrancy and reachable self-destruct operations, reporting them in the source code context. This tool is developed by TrailOfBits [92].
@@
Mythril detects ten types of vulnerabilities, namely, delegatecall, overflow, reentrancy, Use of potentially unsafe/manipulable instructions, Reachable external, Reachable self-destruct instructions, Uninitialized memory usage, Uninitialized storage usage, Enable INVALID instruction detection, and Unused internal transaction return values [93].
5.4.2 Execution Example
@@
For our examination of Manticore, we used the TokenSaleChallenge.sol [85], the same one used to test Oyente. This contract represents a simplified model for trading tokens, which can be seen as stand-ins for fractional tokens of an NFT.
@@
The tool is extremely slow as it takes more than one hour to perform the test; it takes less than a minute for the other tools using the same smart contract and machine. The analysis output is a folder containing a summary report, traces, and analysis for each test case. Manticore generated 51 test cases but failed to detect the smart contact’s reentrancy vulnerability, as shown in Figure 21.
5.4.3 Pros and Cons
@@
Manticore offers a diverse set of command-line tools along with scriptable Python APIs, suitable for a variety of use cases and allowing for in-depth analysis procedures. It covers a broad spectrum of known vulnerabilities, making it a valuable asset for developers seeking to secure their smart contracts.
@@
However, the detailed nature of Manticore’s analysis comes with a trade-off in terms of performance. The tool can take significantly longer to complete its analysis than others, sometimes leading to timeouts. This is often due to its comprehensive symbolic execution approach, which, while powerful, is also resource-intensive. Additionally, Manticore requires considerable memory, which could be a limiting factor for users with constrained hardware resources.
5.5 Echidna
5.5.1 Description